OpenStudy (anonymous):

What the lenght of.an area of.20cm.

OpenStudy (anonymous):

what shape?

OpenStudy (anonymous):

I'll just assume a square, I suppose. A=s^2 for squares, where s is the length of a side. Here, A=20, so 20=s^2 sqrt(20)=s sqrt(4*5)=s 2sqrt(5)=s